OpenStudy (anonymous):

How do I simplify: 3/2m^2n - 1/m^2n^3 + 4/5mn?

OpenStudy (anonymous):

3 1 4 ------ - _____ + _____ 2m^2n m^2n^3 5mn

OpenStudy (anonymous):

So I find for the lowest common denominator: 2m^2n = 2 * m * m * n m^2n^3 = m*m*n*n*n*n 5mn = 5*m*n LCD = 5*2*m*m*n or 10....

OpenStudy (anonymous):

how do i find the lowest common denominator?